1.) Using a Riemann sum with n = 6 subintervals, find the overestimate (i.e. upper Riemann sum) of the area of the region bounded above by the function f(x)= 2-3^x+1 and below by the x-axis on the interval [0,3]. You may give your answer in exact form or in decimal form correct to two decimal places.

2.)On a typical day, a person should consume calories at the rate of c(t) = 50 + 24(root)t-2t^2 calories per hour, where t is the number of hours past midnight. How many calories is consumed between 4 A.M. and 9 A.M.? Please state the units on your answer
Answer #1

Jeg ore to da (641) f(x) = 2-3 [a,b]= [0,3] 0=6 - AN - 3-0 6 2 3 3. 2 left Remans suns = Here & is a decreasing function on [